# Radio over Fiber with Cascaded Structure: Algorithm for Uplink Positioning

### v1

@@ VI-B Validation of CRLB with assumed fiber channels
Figure 8: The magnitude of the frequency response for the frequency-flat fiber and the frequency-selective fiber.
@@
Figure 9: RMSE and CRLB of $\hat{r}$, with frequency flat and frequency selective fibers.
@@
Figure 10: RMSE and CRLB of $\hat{\tau}$, with frequency flat and frequency selective fibers.
@@
To validate the CRLB calculated, extensive Monte-Carlo simulations were performed and based on two artificial channels shown in Fig. 8, corresponding to a flat fiber (red curve) and a frequency selective fiber (blue curve). Each channel contains the same amount of energy, i.e., $\sum_{k=0}^{K-1}\lvert H_{k}\rvert^{2}=\mathcal{E}$. It should be emphasized that the shape of the spectrum of a frequency selective fiber does not impose any deviation in the conclusions of this paper. The corresponding root mean squared error (RMSE) of target parameters was calculated and plotted in Figs. 9, and  10.
@@
The comparison between the CRLB w.r.t. $\hat{r}$ with a flat fiber and that with a non-flat fiber indicates the estimation accuracy is enhanced with a frequency selective fiber, which proves the fact that the system can extract additional information from the spectrum of the received signal. Another observation from Fig. 9 is a flat bound with a flat fiber. This understanding can be acquired from (32i), where the CU estimates the noise variance and determines its quantitative relationship with the noise variance $\sigma^{2}$ at each PA. Therefore, this bound is independent of the exact value of $\sigma^{2}$. The ML estimator achieves an RMSE equal to the CRLB, indicating that it attains optimal performance. Similar conclusions can be drawn from Fig 10. It is worth noting that the gap between the RMSE of certain estimates and the corresponding CRLB does not necessarily indicate suboptimality, as the simulation grids are inherently constrained.
@@
Figure 11: Floorplan of an indoor scenario with three RoFs, a UE trajectory, and the estimated UE trajectory with 1G Hz bandwidth.
Figure 12: Floorplan of an indoor scenario with three RoFs, a UE trajectory, and the estimated UE trajectory with 10G Hz bandwidth.
VII Use case: Indoor positioning
@@
A pertinent use case for the proposed estimation framework on the considered RoF system is indoor positioning. This low-cost design allows the operator to densely deploy RoFs for extreme data rate and good coverage. Here, we assume a UE moves along a trajectory in an indoor scenario illustrated in Fig. 2, where three RoFs are serving the UE simultaneously. The UE is assumed to have an isotropic antenna, while each RU is equipped with a phased array. To ensure fairness and avoid potential biases, no additional assumptions about the antennas are made. Instead, it is simply assumed that the UE is consistently served by three RUs located at three RoFs.
The height of the UE is assumed to be consistently 1.5 meters below the ceiling at all positions, with a uniform spacing of 1 meter between different RUs and between different RoFs. According to (1), the distance between the UE and the entry RU is written as
$d_{m,r}=\sqrt{(P_{x}-r)^{2}+(P_{y}-m)^{2}+1.5^{2}}.$ (38)
Let the vector $\boldsymbol{d}\triangleq[d_{1,r_{1}},d_{2,r_{2}},d_{3,r_{3}}]^{\mathrm{T}}$ store the distances between the UE and three entry RUs. The time-of-flight from the UE to the entry RU is interwoven with a clock offset as defined in (7), which can be viewed as a bias. The position of the UE $P_{x},P_{y}$ is estimated through the estimated $\boldsymbol{\tau}=(1/c)\boldsymbol{d}+\mathbf{1}_{3}\delta_{t}$ at three RoFs and can be written as
$\begin{bmatrix}\hat{P_{x}}\\ \hat{P_{y}}\end{bmatrix}=\mathop{\min}_{P_{x},P_{y},\delta_{\text{t}}}\lvert\lvert\boldsymbol{\tau}-\frac{\boldsymbol{d}}{c}-\mathbf{1}_{3}\delta_{\text{t}}\rvert\rvert^{2},$ (39)
$\begin{bmatrix}\hat{P_{x}}\\ \hat{P_{y}}\end{bmatrix}=\mathop{\min}_{P_{x},P_{y}}\lvert\lvert\boldsymbol{R}-\mathbf{1}_{3}(\mathbf{1}_{3}^{\mathrm{T}}\mathbf{1}_{3})^{-1}\mathbf{1}_{3}^{\mathrm{T}}\boldsymbol{R}\rvert\rvert^{2}.$ (41)
Simulations were conducted using bandwidths of 1 GHz and 10 GHz, corresponding to delay resolutions of 0.3 meters and 0.03 meters, respectively. Figs. 11 and  12 show the estimated UE trajectories with different bandwidths. The RMSE of the estimated positions, obtained through Monte Carlo simulations, is 0.04 meters for 10 GHz and 0.5 meters for 1 GHz, both closely approaching their respective delay resolutions.
@@
4  (23);
5  (25);
6  $c\leftarrow$ (27);
@@
10  $c_{\text{gbest}}\leftarrow 0$$v\leftarrow 0$;
@@
15        $\boldsymbol{\theta}(i,j)\leftarrow(\boldsymbol{\theta}_{\text{max}}(j)-\boldsymbol{\theta}_{\text{min}}(j))*\text{rand(1)}+\boldsymbol{\theta}_{\text{min}}(j)$;
@@ 20        $c_{\text{gbest}}\leftarrow\boldsymbol{c}(i)$;
22           $\boldsymbol{\theta}_{\text{gbest}}(j)\leftarrow\boldsymbol{\theta}(i,j)$;
23
@@ 22           $\boldsymbol{\theta}_{\text{gbest}}(j)\leftarrow\boldsymbol{\theta}(i,j)$;
30     $v\leftarrow w_{\text{ine}}*v+w_{1}*\text{rand(p,n)}*(\theta_{\text{pbest}}-\theta)+w_{2}*\text{rand(p,n)}*(\theta_{\text{gbest}}-\theta)$;
@@ 53              $\boldsymbol{\theta}_{\text{gbest}}(k)\leftarrow\boldsymbol{\theta}(j,k)$;
54
@@
57    $w_{\text{ine}}\leftarrow w_{\text{ine}}*0.7$;
58
59 return $c_{\text{gbest}},\boldsymbol{\theta}_{\text{gbest}}$;
60
61
@@
VIII Conclusion
This paper studied a novel, low-cost, and easily deployable sub-THz RoF system implementation that combines PMFs and RUs in a cascaded structure. An UL signal model of this system was developed by considering two regimes for PAs, linear and non-linear. For our analysis, we considered two type of fibers: flat fibers and frequency selective fibers, whose effects were included in the signal model. Based on this model, an ML framework and a NLS framework were developed to estimate the propagation distance along the RoF and the TOA of signals transmitted from a UE. The CRLB was derived as a benchmark of the variance of the proposed ML estimator.
Monte-Carlo simulations were performed to assess the performance of the proposed estimator. Our results demonstrate that good performance of the proposed estimators can be achieved under the effects of cascaded PMFs and RUs. The ML estimator achieves superior performance for a frequency-selective fiber compared to with a frequency-flat fiber. Finally, our proposed algorithms show high accuracy in positioning UEs in an indoor scenario.
A future direction could be estimating how the second moment of the variance is affected by the non-linearities and including that into the estimation criterion. Another possible direction is to replace the third-order polynomial model of PA non-linearity with a more accurate one.

### v2

@@ VI-B Validation of CRLB with assumed fiber channels
Figure 9: The magnitude of the frequency response for the frequency flat fiber and the frequency-selective fibers.
@@
Figure 10: RMSE and CRLB of $\hat{r}$, with frequency flat and frequency-selective fibers.
@@
Figure 11: RMSE and CRLB of $\hat{\tau}$, with frequency flat and frequency-selective fibers.
@@
To validate the CRLB calculated, extensive Monte-Carlo simulations were performed and based on two artificial channels shown in Fig. 9, corresponding to a flat fiber (red curve) and a frequency-selective fiber (blue curve). Each channel contains the same amount of energy, i.e., $\sum_{k=0}^{K-1}\lvert H_{k}\rvert^{2}=\mathcal{E}$. It should be emphasized that the shape of the spectrum of a frequency-selective fiber does not impose any deviation in the conclusions of this paper. The corresponding root mean squared error (RMSE) of target parameters was calculated and plotted in Figs. 10, and  11.
@@
The comparison between the CRLB w.r.t. $\hat{r}$ with a flat fiber and that with a non-flat fiber indicates that the estimation accuracy is enhanced with a frequency-selective fiber, which proves the fact that the system can extract additional information from the spectrum of the received signal. Another observation from Fig. 10 is a flat bound with a flat fiber. This understanding can be acquired from (35i), where the CU estimates the noise variance and determines its quantitative relationship with the noise variance $\sigma^{2}$ at each PA. Therefore, this bound is independent of the exact value of $\sigma^{2}$. The ML estimator achieves an RMSE equal to the CRLB, indicating that it attains optimal performance. Similar conclusions can be drawn from Fig 11. It is worth noting that the gap between the RMSE of certain estimates and the corresponding CRLB does not necessarily indicate suboptimality, as the simulation grids are inherently constrained.
@@
Figure 12: Floorplan of an indoor scenario with three RoF stripes, a UE trajectory, and the estimated UE trajectory with 1 GHz and 10 GHz bandwidth.
Figure 13: Comparison of RU deployments in parallel and star topologies: (a) parallel-topology network, (b) GDOP distribution for the parallel topology, (c) star-topology network, and (d) GDOP distribution for the star topology.
VI-C Use case: Indoor positioning
@@
A pertinent use case for the proposed estimation framework on the considered RoF system is indoor positioning. This low-cost design allows the operator to densely deploy RoF stripes for extreme data rate and good coverage. Here, we assume a UE moves along a trajectory in an indoor scenario illustrated in Fig. 2, where three RoF stripes are serving the UE simultaneously. To ensure fairness and avoid potential biases, no additional assumptions about the antennas are made. Instead, it is simply assumed that the UE is consistently served by three RUs located at three RoF stripes. The height of the UE is assumed to be consistently 1.5 meters below the ceiling at all positions, with a uniform spacing of 1 meter between adjacent units in stripes and a spacing of 0.5 meters between adjacent RoF stripes.
The simulations were carried out using bandwidths of 1 GHz and 10 GHz, corresponding to delay resolutions of 0.3 meters and 0.03 meters, respectively. Fig 12 shows the estimated UE trajectories with different bandwidths. The RMSE of the estimated positions, obtained through Monte Carlo simulations, is 0.04 meters for 10 GHz and 0.5 meters for 1 GHz, both closely approaching their respective delay resolutions. The GDOP quantifies the influence of the geometry on the positioning performance. The proposed positioning algorithm is based on ToA with bias, whose GDOP can be calculated according to (5)-(6). Stripe deployments can follow different network topologies, such as the parallel topology, which is the focus of this paper, or a star topology (see Figs. 13(a) and 13(c)). For consistency, both topologies are deployed to cover the same service area; however, one should notice that the star topology deploys fewer RUs than the parallel topology. Figs. 13(b) and 13(d) illustrate the GDOP distribution for both topologies. In Figs. 13(b) and  13(d), the high GDOP in certain positions is mainly attributed to the alignment of two RUs along a line, which results in an unfavorable geometry and reduced positioning performance. Additionally, the star network topology exhibits greater geometric diversity than the parallel network topology. A topic for future research can be to investigate optimal RU deployments by jointly considering the positioning and communication performance.
@@
4  $x_{n}=\frac{1}{K}\sum_{k=0}^{K-1}e^{j2\pi\frac{n}{N}k}Ae^{-j2\pi f_{k}\tau}s_{k};$
5 $c\leftarrow$ $\bigg\lvert\bigg\lvert\boldsymbol{y}^{(r)}-f^{r}\left(G\left(x_{n}+\lambda x_{n}\left\lvert x_{n}\right\rvert^{2}\right)\right)\bigg\rvert\bigg\rvert^{2}\;$
@@ 4  $x_{n}=\frac{1}{K}\sum_{k=0}^{K-1}e^{j2\pi\frac{n}{N}k}Ae^{-j2\pi f_{k}\tau}s_{k};$
  $c_{\text{gbest}}\leftarrow 0$$v\leftarrow 0$ ;
  // Initialize global cost
@@
        $\boldsymbol{\theta}(i,j)\leftarrow(\boldsymbol{\theta}_{\text{max}}(j)-\boldsymbol{\theta}_{\text{min}}(j))*\text{rand(1)}+\boldsymbol{\theta}_{\text{min}}(j)$;
        // Set the initial values
@@ 17        $c_{\text{gbest}}\leftarrow\boldsymbol{c}(i)$;
           $\boldsymbol{\theta}_{\text{gbest}}(j)\leftarrow\boldsymbol{\theta}(i,j)$;
           // Iterate the global best cost
19
20
21
22
@@
     $v\leftarrow w_{\text{ine}}*v+w_{1}*\text{rand(p,n)}*(\theta_{\text{pbest}}-\theta)+w_{2}*\text{rand(p,n)}*(\theta_{\text{gbest}}-\theta)$;
     // Move particles
@@ 48              $\boldsymbol{\theta}_{\text{gbest}}(k)\leftarrow\boldsymbol{\theta}(j,k)$;
49
50
51
52    $w_{\text{ine}}\leftarrow w_{\text{ine}}*0.7$;
53
54 return $c_{\text{gbest}},\boldsymbol{\theta}_{\text{gbest}}$;
@@
VII Conclusion
This paper studied a novel, low-cost, and easily deployable sub-THz RoF system implementation that combines PMFs and RUs in a cascaded structure. A UL signal model of this system was developed by considering two regimes for PAs, linear and non-linear. For our analysis, we considered two types of fibers: flat fibers and frequency-selective fibers, whose effects were included in the signal model. Based on this model, an ML framework and an NLS framework were developed to estimate the propagation distance along the RoF and the TOA of signals transmitted from a UE. The CRLB was derived as a benchmark of the variance of the proposed ML estimator.
Monte-Carlo simulations were performed to assess the performance of the proposed estimator. Our results demonstrate that good performance of the proposed estimators can be achieved under the effects of cascaded PMFs and RUs. The ML estimator achieves superior performance for a frequency-selective fiber compared to a frequency-flat fiber. Finally, our proposed algorithms show high accuracy in positioning UEs in an indoor scenario.
A future direction could be estimating how the second moment of the variance is affected by the non-linearities and including that in the estimation criterion. Another possible direction is to replace the third-order polynomial model of PA non-linearity with a more accurate one. Also, considering the antenna radiation patterns, such as a highly directional antenna, and handling the UE’s orientation, enhances the research.
@@ Acknowledgment
In this appendix, we provide the detailed derivations of the CRLB shown in Section V. Specifically, we write the respective derivations of the CRLB for both the flat-fiber and frequency-selective fiber assumptions. Note that both CRLBs are obtained from (12). The two types of fiber show different characteristics. Hence, we present the CRLB for each case separately. We define $\boldsymbol{\Omega}_{k}\triangleq\phi+r\psi_{k}-2\pi f_{k}\tau$.
CRLB for flat fiber
Partial derivatives in (33) can be calculated as
$\left[\frac{\partial\boldsymbol{\mu(\theta)}}{\partial\lvert A\rvert}\right]_{k}=$ $Ge^{j(\boldsymbol{\Omega}_{k})}s_{k},$ (40a)
$\left[\frac{\partial\boldsymbol{\mu(\theta)}}{\partial\phi}\right]_{k}=$ $jG\lvert A\rvert s_{k}e^{j(\boldsymbol{\Omega}_{k})},$ (40b)
$\left[\frac{\partial\boldsymbol{\mu(\theta)}}{\partial\tau}\right]_{k}=$ $-j2\pi f_{k}G\lvert A\rvert s_{k}e^{j(\boldsymbol{\Omega}_{k})},$ (40c)
$\left[\frac{\partial\boldsymbol{\mu(\theta)}}{\partial r}\right]_{k}=$ $G\lvert A\rvert e^{j(\boldsymbol{\Omega}_{k})}s_{k}\left(j\psi_{k}\right).$ (40d)
The derivations of CRLB are as follows.
$\left[\boldsymbol{I}(\theta)\right]_{\lvert A\rvert,\lvert A\rvert}$= $2\Re\left\{\!\frac{G^{(2r+2)}}{(r+1)\sigma^{2}}\!\sum_{k=0}^{K-1}\lvert H_{k}\rvert^{2r}\!\lvert s_{k}\rvert^{2}\!e^{-j(\boldsymbol{\Omega}_{k})}\!e^{j(\boldsymbol{\Omega}_{k})}\!\right\},$
$=2\frac{G^{2}}{(r+1)\sigma^{2}}\sum_{k=0}^{K-1}\lvert s_{k}\rvert^{2},$ (41a)
$\left[\boldsymbol{I}(\theta)\right]_{\lvert A\rvert,\phi}=$ $2\Re\left\{\frac{j\lvert A\rvert G^{2}}{(r+1)\sigma^{2}}\sum_{k=0}^{K-1}\lvert s_{k}\rvert^{2}e^{-j(\boldsymbol{\Omega}_{k})}e^{j(\boldsymbol{\Omega}_{k})}\right\},$
$=$ $0,$ (41b)
$\left[\boldsymbol{I}(\theta)\right]_{\lvert A\rvert,\tau}=$ $2\Re\left\{-\frac{j2\pi G^{2}\lvert A\rvert}{(r+1)\sigma^{2}}\sum_{k=0}^{K-1}f_{k}\lvert s_{k}\rvert^{2}e^{-j(\boldsymbol{\Omega}_{k})}e^{j(\boldsymbol{\Omega}_{k})}\right\},$
$=$ $0,$ (41c)
$\left[\boldsymbol{I}(\theta)\right]_{\lvert A\rvert,r}=$ $2\Re\left\{\frac{G^{2}\lvert A\rvert}{(r+1)\sigma^{2}}\sum_{k=0}^{K-1}\lvert s_{k}\rvert^{2}\left(j\psi_{k}\right)\right\},$
$=$ $0,$ (41d)
$\left[\boldsymbol{I}(\theta)\right]_{\phi,\phi}=$ $2\Re\left\{\frac{-j^{2}\lvert A\rvert^{2}G^{2}}{(r+1)\sigma^{2}}\sum_{k=0}^{K-1}\lvert s_{k}\rvert^{2}e^{-j(\boldsymbol{\Omega}_{k})}e^{j(\boldsymbol{\Omega}_{k})}\right\},$
$=$ $2\frac{G^{2}\lvert A\rvert^{2}}{(r+1)\sigma^{2}}\sum_{k=0}^{K-1}\lvert s_{k}\rvert^{2},$ (41e)
$\left[\boldsymbol{I}(\theta)\right]_{\phi,\tau}=$ $2\Re\left\{\frac{j^{2}2\pi\lvert A\rvert^{2}G^{2}}{(r+1)\sigma^{2}}\sum_{k=0}^{K-1}f_{k}\lvert s_{k}\rvert^{2}e^{-j(\boldsymbol{\Omega}_{k})}e^{j(\boldsymbol{\Omega}_{k})}\right\},$
$=$ $-\frac{4\pi G^{2}\lvert A\rvert^{2}}{(r+1)\sigma^{2}}\sum_{k=0}^{K-1}f_{k}\lvert s_{k}\rvert^{2},$ (41f)
$\left[\boldsymbol{I}(\theta)\right]_{\phi,r}=$ $2\Re\left\{\frac{-j\lvert A\rvert^{2}G^{2}}{(r+1)\sigma^{2}}\sum_{k=0}^{K-1}\lvert s_{k}\rvert^{2}\left(j\psi_{k}\right)\right\},$
$=$ $2\frac{G^{2}\lvert A\rvert^{2}}{(r+1)\sigma^{2}}\sum_{k=0}^{K-1}\lvert s_{k}\rvert^{2}\psi_{k},$ (41g)
$\left[\boldsymbol{I}(\theta)\right]_{\tau,\tau}=$ $2\Re\left\{\frac{-j^{2}4\pi^{2}\lvert A\rvert^{2}G^{2}}{(r+1)\sigma^{2}}\sum_{k=0}^{K-1}f_{k}^{2}\lvert s_{k}\rvert^{2}e^{-j(\boldsymbol{\Omega}_{k})}e^{j(\boldsymbol{\Omega}_{k})}\right\},$
$=$ $\frac{8\pi^{2}G^{2}\lvert A\rvert^{2}}{(r+1)\sigma^{2}}\sum_{k=0}^{K-1}f_{k}^{2}\lvert s_{k}\rvert^{2},$ (41h)
$\left[\boldsymbol{I}(\theta)\right]_{r,r}=$ $K\left\{\frac{\sigma^{2}}{(r+1)\sigma^{2}}\frac{\sigma^{2}}{(r+1)\sigma^{2}}\right\}$
$+2\Re\left\{\frac{\lvert A\rvert^{2}G^{2}}{(r+1)\sigma^{2}}\sum_{k=0}^{K-1}\lvert s_{k}\rvert^{2}\left(-j\psi_{k}\right)\left(j\psi_{k}\right)\right\},$
$=$ $\frac{K}{(r+1)^{2}}+2\frac{G^{2}\lvert A\rvert^{2}}{(r+1)\sigma^{2}}\sum_{k=0}^{K-1}\lvert s_{k}\rvert^{2}\psi_{k}^{2},$ (41i)
$\left[\boldsymbol{I}(\theta)\right]_{\tau,r}=$ $2\Re\left\{\frac{j2\pi\lvert A\rvert^{2}G^{2}}{(r+1)\sigma^{2}}\sum_{k=0}^{K-1}f_{k}\lvert s_{k}\rvert^{2}\left(j\psi_{k}\right)\right\},$
$=$ $-\frac{4\pi G^{2}\lvert A\rvert^{2}}{(r+1)\sigma^{2}}\sum_{k=0}^{K-1}\lvert s_{k}\rvert^{2}f_{k}\psi_{k}.$ (41j)
CRLB for frequency-selective fiber
One can derive the first-order derivatives in (36) as
$\times\left(\!\ln\!{\sqrt{b_{k}}}\!+\!j\psi_{k}\!\right)\!.$ (42d)
The derivations of CRLB as as follows.
$\left[\boldsymbol{I}(\theta)\right]_{\lvert A\rvert,\lvert A\rvert}$= $2\Re\left\{\!G^{2}\!\sum_{k=0}^{K-1}\!\frac{(b_{k}\!-\!1)b_{k}^{r}\lvert s_{k}\rvert^{2}\!e^{-j(\boldsymbol{\Omega}_{k})}e^{j(\boldsymbol{\Omega}_{k})}}{\sigma^{2}(b_{k}^{r+1}-1)}\!\right\},$
$=2G^{2}\sum_{k=0}^{K-1}\frac{(b_{k}-1)b_{k}^{r}\lvert s_{k}\rvert^{2}}{\sigma^{2}(b_{k}^{r+1}-1)},$ (43a)
$\left[\boldsymbol{I}(\theta)\right]_{\lvert A\rvert,\phi}=$ $2\Re\left\{j\lvert A\rvert G^{2}\sum_{k=0}^{K-1}\frac{(b_{k}-1)b_{k}^{r}\lvert s_{k}\rvert^{2}e^{-j(\boldsymbol{\Omega}_{k})}e^{j(\boldsymbol{\Omega}_{k})}}{\sigma^{2}(b_{k}^{r+1}-1)}\right\},$
$=$ $0,$ (43b)
$\left[\boldsymbol{I}(\theta)\right]_{\lvert A\rvert,\tau}=$ $2\Re\left\{\!-j2\pi G^{2}\lvert A\rvert\!\sum_{k=0}^{K-1}\!\frac{(b_{k}-1)f_{k}b_{k}^{r}\lvert s_{k}\rvert^{2}e^{-j(\boldsymbol{\Omega}_{k})}e^{j(\boldsymbol{\Omega}_{k})}}{\sigma^{2}(b_{k}^{r+1}-1)}\!\right\},$
$=$ $0,$ (43c)
$\left[\boldsymbol{I}(\theta)\right]_{\lvert A\rvert,r}=$ $2\Re\left\{G^{2}\lvert A\rvert\sum_{k=0}^{K-1}\frac{(b_{k}-1)b_{k}^{r}\lvert s_{k}\rvert^{2}\left(\ln{\lvert GH_{k}\rvert}+j\psi_{k}\right)}{\sigma^{2}(b_{k}^{r+1}-1)}\right\},$
$=$ $2G^{2}\lvert A\rvert\sum_{k=0}^{K-1}\frac{(b_{k}-1)b_{k}^{r}\lvert s_{k}\rvert^{2}\ln{\sqrt{b_{k}}}}{\sigma^{2}(b_{k}^{r+1}-1)},$ (43d)
$\left[\boldsymbol{I}(\theta)\right]_{\phi,\phi}=$ $2\Re\left\{\lvert A\rvert^{2}G^{2}\sum_{k=0}^{K-1}\frac{(b_{k}-1)b_{k}^{r}\lvert s_{k}\rvert^{2}e^{-j(\boldsymbol{\Omega}_{k})}e^{j(\boldsymbol{\Omega}_{k})}}{\sigma^{2}(b_{k}^{r+1}-1)}\right\},$
$=$ $2G^{2}\lvert A\rvert^{2}\sum_{k=0}^{K-1}\frac{(b_{k}-1)b_{k}^{r}\lvert s_{k}\rvert^{2}}{\sigma^{2}(b_{k}^{r+1}-1)},$ (43e)
$\left[\boldsymbol{I}(\theta)\right]_{\phi,\tau}=$ $2\Re\left\{\!-2\pi\lvert A\rvert^{2}G^{2}\!\sum_{k=0}^{K-1}\!\frac{(b_{k}-1)f_{k}b_{k}^{r}\lvert s_{k}\rvert^{2}e^{-j(\boldsymbol{\Omega}_{k})}e^{j(\boldsymbol{\Omega}_{k})}}{\sigma^{2}(b_{k}^{r+1}-1)}\!\right\},$
$=$ $-4\pi G^{2}\lvert A\rvert^{2}\sum_{k=0}^{K-1}\frac{(b_{k}-1)f_{k}b_{k}^{r}\lvert s_{k}\rvert^{2}}{\sigma^{2}(b_{k}^{r+1}-1)},$ (43f)
$\left[\boldsymbol{I}(\theta)\right]_{\phi,r}=$ $2\Re\left\{\!-j\lvert A\rvert^{2}G^{2}\!\sum_{k=0}^{K-1}\!\frac{(b_{k}-1)b_{k}^{r}\lvert s_{k}\rvert^{2}\left(\ln{\lvert GH_{k}\rvert}+j\psi_{k}\right)}{\sigma^{2}(b_{k}^{r+1}-1)}\!\right\},$
$=$ $2G^{2}\lvert A\rvert^{2}\sum_{k=0}^{K-1}\frac{(b_{k}-1)b_{k}^{r}\lvert s_{k}\rvert^{2}\psi_{k}}{\sigma^{2}(b_{k}^{r+1}-1)},$ (43g)
$\left[\boldsymbol{I}(\theta)\right]_{\tau,\tau}=$ $2\Re\left\{\!4\pi^{2}\lvert A\rvert^{2}G^{2}\!\sum_{k=0}^{K-1}\!\frac{(b_{k}-1)f_{k}^{2}b_{k}^{r}\lvert s_{k}\rvert^{2}e^{-j(\boldsymbol{\Omega}_{k})}e^{j(\boldsymbol{\Omega}_{k})}}{\sigma^{2}(b_{k}^{r+1}-1)}\!\right\},$
$=$ $8\pi^{2}G^{2}\lvert A\rvert^{2}\sum_{k=0}^{K-1}\frac{(b_{k}-1)f_{k}^{2}b_{k}^{r}\lvert s_{k}\rvert^{2}}{\sigma^{2}(b_{k}^{r+1}-1)},$ (43h)
$\left[\boldsymbol{I}(\theta)\right]_{r,r}=$ $\mathrm{tr}\left\{\boldsymbol{C}(\boldsymbol{\theta})^{-1}\frac{\partial\boldsymbol{C}(\boldsymbol{\theta})}{\partial r}\boldsymbol{C}(\boldsymbol{\theta})^{-1}\frac{\partial\boldsymbol{C}(\boldsymbol{\theta})}{\partial r}\right\}+2\Re\left\{\lvert A\rvert^{2}G^{2}\sum_{k=0}^{K-1}\frac{(b_{k}-1)\lvert s_{k}\rvert^{2}b_{k}^{r}\left(\ln{\lvert GH_{k}\rvert}-j\psi_{k}\right)\left(\ln{\lvert GH_{k}\rvert}+j\psi_{k}\right)}{\sigma^{2}(b_{k}^{r+1}-1)}\right\},$
$=$ $\sum_{k=0}^{K-1}\left(\frac{b_{k}^{r}\ln{b_{k}}}{1-b_{k}^{r+1}}\right)^{2}+2G^{2}\lvert A\rvert^{2}\sum_{k=0}^{K-1}\frac{\lvert s_{k}\rvert^{2}(b_{k}-1)b_{k}^{r}\left((\ln{\sqrt{b_{k}}})^{2}+\psi_{k}^{2}\right)}{\sigma^{2}(b_{k}^{r+1}-1)},$ (43i)
$\left[\boldsymbol{I}(\theta)\right]_{\tau,r}=$ $2\Re\left\{j2\pi\lvert A\rvert^{2}G^{2}\sum_{k=0}^{K-1}\frac{(b_{k}-1)f_{k}b_{k}^{r}\lvert s_{k}\rvert^{2}\left(\ln{\lvert GH_{k}\rvert}+j\psi_{k}\right)}{\sigma^{2}(b_{k}^{r+1}-1)}\right\},$
$=$ $-4\pi G^{2}\lvert A\rvert^{2}\sum_{k=0}^{K-1}\frac{(b_{k}-1)b_{k}^{r}\lvert s_{k}\rvert^{2}f_{k}\psi_{k}}{\sigma^{2}(b_{k}^{r+1}-1)}.$ (43j)
[Uncaptioned image] Dexin Kong (Graduate Student Member, IEEE) received his B.Eng. degree in communication engineering from Nanjing University of Posts and Telecommunications (NJUPT), Nanjing, China, in 2021, and M.Sc. degree in communication systems from Lund University, Lund, Sweden, in 2023. He is currently pursuing a Ph.D. degree with the Department of Electrical Engineering at Linköping University, Linköping, Sweden. His research interests include wireless communications, sub-terahertz communications, and radio-over-fiber technologies.
[Uncaptioned image] Diana Pamela Moya Osorio (M’16, SM’23) is currently Associate Professor at the Communication Systems Division, Department of Electrical Engineering, Linköping University, Sweden, and an ELLIIT recruited faculty. Previously, she was Senior Research Fellow and Adjunct Professor at the Centre for Wireless Communications, University of Oulu, Finland. She received the B.Sc. degree in electronics and telecommunications engineering from the Armed Forces University, Ecuador, in 2008, and the M.Sc. and D.Sc. degrees in electrical engineering with emphasis on telecommunications and telematics from the University of Campinas, Brazil, in 2011 and 2015, respectively. From 2015 to 2022, she was an Assistant Professor with the Department of Electrical Engineering, Federal University of São Carlos, Brazil. From 2020 to 2023, she was also a Postdoctoral Researcher for the Academy of Finland. She has served as TPC and reviewer for several journals and conferences. Currently, she is Associate Editor of IEEE Communication Letters and IEEE Transactions on Information Forensics & Security. She also serves as working group leader for Trustworthy 6G at the Cost Action 6G-PHYSEC. Her research interests include wireless communications and sensing in general, signal processing for wireless communications and radar systems, physical layer security and trustworthy systems, and integrated sensing and communications.
[Uncaptioned image] Erik G. Larsson (Fellow, IEEE) received the Ph.D. degree from Uppsala University, Uppsala, Sweden, in 2002. He is currently Professor of Communication Systems at Linköping University (LiU) in Linköping, Sweden. He was with the KTH Royal Institute of Technology in Stockholm, Sweden, the George Washington University, USA, the University of Florida, USA, and Ericsson Research, Sweden. His main professional interests are within wireless communications, signal processing, network science, and decentralized machine learning. He co-authored Space-Time Block Coding for Wireless Communications (Cambridge University Press, 2003) and Fundamentals of Massive MIMO (Cambridge University Press, 2016). He served as chair of the IEEE Signal Processing Society SPCOM technical committee (2015–2016), chair of the IEEE Wireless Communications Letters steering committee (2014–2015), member of the IEEE Transactions on Wireless Communications steering committee (2019-2022), General and Technical Chair of the Asilomar SSC conference (2015, 2012), technical co-chair of the IEEE Communication Theory Workshop (2019), and member of the IEEE Signal Processing Society Awards Board (2017–2019). He was Associate Editor for, among others, the IEEE Transactions on Communications (2010-2014), the IEEE Transactions on Signal Processing (2006-2010), and the IEEE Signal Processing Magazine (2018-2022). He received the IEEE Signal Processing Magazine Best Column Award twice, in 2012 and 2014, the IEEE ComSoc Stephen O. Rice Prize in Communications Theory in 2015, the IEEE ComSoc Leonard G. Abraham Prize in 2017, the IEEE ComSoc Best Tutorial Paper Award in 2018, the IEEE ComSoc Fred W. Ellersick Prize in 2019, and the IEEE SPS Donald G. Fink Overview Paper Award in 2023. He is a Fellow of EURASIP, a member of the Swedish Royal Academy of Sciences (KVA), and Highly Cited according to ISI Web of Science.